An electron-positron pair is produced by a 2.30 MeV photon. What is the kinetic energy of the positron if the kinetic energy of the electron is 0.958 MeV? Use the following Joules-to-electron-Volts conversion 1eV = 1.602 × 10-19 J. The rest mass of an electron is 9.11 × 10-31 kg
Answer #1

KEpair = KEelectron + KEpositron  (Basic concept of energy conservation)

2.30 = 0.958 + KEpositron (in MeV)

KEpositron = 1.342 MeV

KEpositron = 1.342 * 106 eV = 1.342 * 106 * (1.602 * 10-19) J

KEpositron = 2.15 * 10-13 J